OpenStudy (anonymous):

I have a Math question, I really need help on this. The sum of the measures of angle x and angle y is 90. If the measures of angle x is 30 less than twice the measure of angle y, what is the measure of angle x? ( Choices ) 20 35 50 65

OpenStudy (anonymous):

It cant be 65 65*2 - 30 = 100 100+65 =165 90=165 it cant be 50 50*2-30=70 70+50=130 90=130 it cant be 35 35*2-30=40 40+35=75 75=90 it cant be 20 20-2-30=10 20+10=30 30=90

OpenStudy (anonymous):

so i have know idea

OpenStudy (mathstudent55):

How do you write as an expression "the sum of the measures of angle x and angle y"? Then let the expression equal 90. That gives one equation. Now write twice y minus 30, and let that equal x. That is your second equation. Now solve the two equations simultaneously, as a system of equations, and find x and y.

OpenStudy (anonymous):

So is it 35?

OpenStudy (mathstudent55):

@megan.brennan made a mistake in working out the problem starting with the solutions. The question is what is x? The choices are possible values of x. The way megan used the choices was by replacing y in the equation x = 2y - 30. That's why none of the answers worked.

OpenStudy (mathstudent55):

One of the choices is definitely correct. How about following what I wrote, writing a system of equations, and solving the system?

OpenStudy (anonymous):

This is my worst subject, I have no idea of what you said...

OpenStudy (anonymous):

Thats not rude at all. @mathstudent55

OpenStudy (anonymous):

you have a cable plan that costs 39$ a month for a basic plan plus one movie channel. your friend has the same basic plan plus two movie channels for 45.50$. what is the basic plan charge that you both pay? @mathstudent55 why your at it

OpenStudy (anonymous):

So it would be $6?

OpenStudy (mathstudent55):

basic + 1 movie channel = $39 basic + 2 movie channels = $45.50 Let b = the cost of basic cable Let m = the cost of a movie channel you have: b + m = 39 your friend has: b + 2m = 45.50 We have a system of equations: b + m = 39 b + 2m = 45.50 Subtract the first equation from the second equation: b + 2m = 45.50 - b + m = 39 -------------------- m = 6.50 m = 6.5 means one movie channel costs $6.50

OpenStudy (anonymous):

So is the answer 60? ( D)

OpenStudy (mathstudent55):

BTW, with the cable problem, you can think of it simply like this, and not even write an equation. The plans cost $39 and $45.50. Your plan is basic & 1 movie channel Your friend's plan is basic and 2 movie channels. What is the difference between your plan and your friend's plan? Just 1 movie channel. He gets an extra movie channel that you don't. How much does he pay more than you do? $45.50 - $39 = $6.50 He pay an extra $6.50. Since the only difference between your plan and his is one movie channel, that means the movie channel costs $6.50.

OpenStudy (mathstudent55):

Now we know in two different ways that the movie channels cost $6.50 each. Now we can continue to answer the original question. What is the cost of basic cable? Using the equations: b + m = 39 b + 6.50 = 39 b = 32.50 Basic cable costs $32.50

OpenStudy (anonymous):

I'm just lost now xD

OpenStudy (mathstudent55):

Without using the equation, you can also find the cost of basic cable. If your cost is $39 for basic and 1 movie channel, and 1 movie channel costs $6.50, then $39 - $6.50 is basic cost. basic cable cost = $39 - $6.50 = $32.50 Basic cable costs $32.50

OpenStudy (mathstudent55):

Ok. Now we should get back to @TimaFilms99 original question about angles because we are still in Tima's post.

OpenStudy (mathstudent55):

@TimaFilms99 I'm going to help you now with your question in a simpler way without equations. Try to follow.

OpenStudy (anonymous):

Yes please xD

OpenStudy (mathstudent55):

You have two angles and they are called x and y. You are told that their measures add to 90 degrees. For example, if angle x were 10 degrees, then angle y would have to be 80 degrees. If angle x were 30 degrees, then angle y would have to be 60 degrees. This is because 80 + 10 = 90 and 30 + 60 = 90 in our examples. Ok so far?

OpenStudy (anonymous):

I will try, I have a D in this class.

OpenStudy (anonymous):

Yea

OpenStudy (mathstudent55):

You are also given a second statement about the measures of angles x and y. You are told that angle x is 30 less than twice y. For example, if angle y were 60, then to find angle x, you need to multiply y by 2 to get 120, then subtract 30 to get 90. You would have x = 60 and y = 90.

OpenStudy (mathstudent55):

Now we need to use the two statements to find out what the measures of x and y are.

OpenStudy (mathstudent55):

Instead of just guessing numbers at random, we can use the choices. We can try each choice and see which one works.

OpenStudy (mathstudent55):

Choice A is x = 20. If x = 20, then y must be 70 since 20 + 70 = 90. Now we have to see if 20 and 70 work in the second statement. x = 2 * 70 - 30 = 140 - 30 = 110 Now we get x = 110, but above x had to be 20, so choice A does not work.

OpenStudy (mathstudent55):

Since choice A did not work, we try choice B. Choice B is x = 35. If x = 35, then y must be 55 since 35 + 55 = 90. Now we have to see if 35 and 55 work in the second statement. x = 2 * 55 - 30 = 110 - 30 = 80 Now we get x = 80, but above x had to be 35, so choice B does not work.

OpenStudy (mathstudent55):

Since choice B did not work, we try choice C. Choice B is x = 50. If x = 50, then y must be 40 since 50 + 40 = 90. Now we have to see if 50 and 40 work in the second statement. x = 2 * 40 - 30 = 80 - 30 = 50 Now we get x = 50, and above we also had x = 50, so this is the answer. The answer is choice C.

OpenStudy (anonymous):

At an all-you-can-eat barbecue fundraiser that you are sponsoring, adults pay 6$ for a dinner and children pay 4$ for a dinner. 212 people attend and raise 1128$. What is the total number of adults and the total number of children attending. (using systems of equations. a. what is a system of equations that you can solve this problem? b.what method would you use and why?

OpenStudy (anonymous):

Sorry I had to go, and I still don't know...